OpenStudy (anonymous):

prove that tan^2θ+1sec^2θ for all θ and use 80 degrees to verify the identity

OpenStudy (anonymous):

what is tan^2 and sec^2 equal to?

OpenStudy (anonymous):

i will make a guess that this is \[\tan^2(\theta)+1=\sec^2(\theta)\]

OpenStudy (anonymous):

take your famous pythagorean identity and divide everything by \(\cos^2(\theta)\)

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

Rewrite the left side in sines and cosines: sin^2 (theta) / cos^2 (theta) + 1 = sin^2 (theta)/cos^2 (theta) + cos^2(theta)/cos^2 (theta) = (sin^2 (theta) + cos^2 (theta)) / cos^2 (theta) = 1 / cos^2 (theta) = sec^2 (theta) At 80 degrees: tan^2(80) + 1 = (5.67128182)^2 + 1 = 33.16343748 sec^2 (80) = (5.758770483)^2 = 33.16343748

OpenStudy (anonymous):

hope you can read it

OpenStudy (anonymous):

do u have a brighter photo u cant really read it

OpenStudy (anonymous):

yea just a sec though

OpenStudy (anonymous):

Rewrite the left side in sines and cosines: sin^2 (theta) / cos^2 (theta) + 1 = sin^2 (theta)/cos^2 (theta) + cos^2(theta)/cos^2 (theta) = (sin^2 (theta) + cos^2 (theta)) / cos^2 (theta) = 1 / cos^2 (theta) = sec^2 (theta) At 80 degrees: tan^2(80) + 1 = (5.67128182)^2 + 1 = 33.16343748 sec^2 (80) = (5.758770483)^2 = 33.16343748 this is the answer i think so here u go fari321

OpenStudy (anonymous):

this is as good as it gets

OpenStudy (anonymous):

then just plug in the angles and they will equal 1

OpenStudy (anonymous):

plug in 80 degrees?

OpenStudy (anonymous):

yes, to verify the identity just say sin(80)^2 + cos(80)^2 =1